• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 21:34:59 on November 24, 1998

    Entry number 6830

    keyword=Beam Time Accounting
    Hall A

    11/24/1998 1600-2400
    Experiment #: E94-010
    Shift Leader: D.S. Brown
    Shift Workers: A. Deur, M. Kuss, H. Ueno, R. DiSalvo

    16001700180019002000210022002300Total Time
    Acceptable Beam
    in Use
    30 50 1 1 30 0 0 0 3.83333333333333
    Beam Available
    but Not in Use
    30 10 0 0 30 0 0 0 1.16666666666667
    Beam Not Available
    or Unacceptable
    0 0 0 0 0 1 1 1 3
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    30 50 1 1 30 0 0 0 3.83333333333333
    Planned Config-
    uration Change
    30 10 0 0 30 1 1 1 4.16666666666667
    Unplanned
    Experiment Down
    0 0 0 0 0 0 0 00
    Hall not
    ready reasons
    Target 0 0 0 0 0 0 0 0 0
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: Optics studies completed during shift; we finshed the 2% and 4% settings. We used remaining beam time after 20:00 to take short He3 run at p0=863.8 MeV, and 2 quasielastic runs on carbon at p0=700 MeV (no collimator and 6msr). The rest of shift will have He3 cell swapped out with H20 cell for calibration and then new He3 cell installed for 3-pass running after Thanksgiving.

    Data Entry by: D.S. Brown